A laboratory technician needs to make a 42-liter batch of a 40% acid solution. How can the laboratory technician combine a batch of an acid solution that is pure acid with another that is 10% to get the desired concentration?
Answer by macston(5194)

.
P=amount of pure acid; T=amount of 10% solution
P+T=42L
T=42L-P
.
(1.00)P+(0.10)T=(0.40)42L
P+(0.10)(42L-P)=16.8L
P+4.2L-0.10P=16.8L
0.90P=12.6L
P=14
ANSWER 1: He should use 14 liters of pure acid.
T=42L-P=42L-14L=28L
ANSWER 2: He should use 28 liters of 10% acid solution.
.
CHECK:
(1.00)P+(0.10)T=(0.40)42L
14L+(0.10)28L=16.8L
14L+2.8L=16.8L
16.8L=16.8L